### v2

# Araida: Analogical Reasoning-Augmented Interactive Data Annotation

Abstract
Human annotation is a time-consuming task that requires a significant amount of effort. To address this issue, interactive data annotation utilizes an annotation model to provide suggestions for humans to approve or correct. However, annotation models trained with limited labeled data are prone to generating incorrect suggestions, leading to extra human correction effort. To tackle this challenge, we propose Araida, an analogical reasoning-based approach that enhances automatic annotation accuracy in the interactive data annotation setting and reduces the need for human corrections. Araida involves an error-aware integration strategy that dynamically coordinates an annotation model and a k-nearest neighbors (KNN) model, giving more importance to KNN’s predictions when predictions from the annotation model are deemed inaccurate. Empirical studies demonstrate that Araida is adaptable to different annotation tasks and models. On average, it reduces human correction labor by 11.02% compared to vanilla interactive data annotation methods.
Araida: Analogical Reasoning-Augmented Interactive Data Annotation
Chen Huang♠♣,  Yiping Jin,  Ilija Ilievski,  Wenqiang Lei♠♣thanks: Correspondence to Wenqiang Lei.,  Jiancheng Lv♠♣ College of Computer Science, Sichuan University, China Engineering Research Center of Machine Learning and Industry Intelligence, Ministry of Education, China NLP Group, Pompeu Fabra University, Spain ISEM, National University of Singapore, Singapore huangc.scu@gmail.com, yiping.jin@upf.edu, wenqianglei@gmail.com
1 Introduction
Data annotation is a challenging task that involves a tradeoff between annotation quality and budget. While some platforms offer a cost-effective solution by relying on ML models to annotate data automatically  [^1] For example, https://aws.amazon.com/sagemaker/groundtruth/., the quality of such annotations is often compromised Wang et al. (2022a). It is particularly true in the limited data annotation scenario where the annotation budget is limited or when unlabeled data are scarce Ringger et al. (2007); Chaudhary et al. (2021); Huang et al. (2024).
Human-machine interactive annotation methods were introduced to reduce annotation effort while maintaining annotation quality Klie et al. (2018, 2020); Le et al. (2021). As illustrated in Fig. 1, these methods introduce an annotation model to suggest labels (model annotations) to human annotators. The annotators accept a suggested label if it is correct. Otherwise, they have to correct the label manually. Compared to manual annotation, interactive annotation requires less human effort because human annotators only have to verify the model annotations instead of coming up with an answer from scratch, leading to potential speedup of the annotation process (Klie et al., 2020).
Figure 1: Comparison between manual annotation and interactive annotation.
Figure 2: Example on span relation annotation. An under-trained annotation model results in more suggestion errors and increases human correction effort. Araida improves the model annotation accuracy via the KNN model and the error-aware integration strategy for dynamical coordination of annotations.
Evidently, the annotation model’s accuracy is crucial because incorrect suggestions require additional human effort to rectify. Existing methods update the annotation model based on previously accepted or corrected data (ground-truth annotation), aiming to reduce human corrections by improving prediction accuracy at each iteration Klie et al. (2020); Wu et al. (2022). However, in the context of limited data annotation, the annotation model lacks sufficient labeled training data to reach a reasonable accuracy and is prone to providing incorrect suggestions Rietz and Maedche (2021). For example, in the span relation annotation example shown in Fig. 2 (blue), the annotation model continues to make mistakes on similar examples ([car, tyre]) even after the human annotator corrects the label ‘[tree, leaf]=>absent=>= >component’. As a result, this leads to more human corrections. Such a problem is crucial for interactive annotation and has been identified by recent work Rietz and Maedche (2021), but it has yet to be addressed.
Inspired by cognitive studies on efficient learning Lake et al. (2017, 2015); Mitchell (2021), finding that the human brain can learn from a few examples because our brain is continuously building analogies during the learning process of concepts to facilitate comprehension, we propose Analogical Reasoning-Augmented Interactive Data Annotation (Araida), which is designed to improve interactive annotation under the limited data annotation setting. Araida provides an annotation reference to the annotation model by retrieving previously human-labeled examples in the proximity of the example in consideration using the k-nearest neighbors (KNN) method. As illustrated in Fig. 2(red), the final suggestion combines the model annotation and the annotation reference provided by KNN via an error-aware integration strategy. This strategy dynamically coordinates the annotation model and KNN, giving more importance to KNN’s prediction if the predicted label from the annotation model is estimated to be inaccurate.
We conduct simulated experiments for the limited data annotation task and estimate the human annotation effort based on the number of human corrections (or the number of suggestion errors) following Hwa (2000) and Kristjansson et al. (2004). We test Araida on different word-level and sentence-level annotation tasks, combining with different annotation models (i.e., classic and LLM-based models). The result shows that Araida consistently improves different annotation models’ accuracy across various tasks. On average, it reduces human corrections by 11.02%. Further analysis attributes this improvement to the few-shot capability of the KNN module and the error-aware integration strategy that effectively synergizes complementary annotations. In summary, our contributions are as follows:
  • Calling attention to the limited data annotation scenario. We highlight the under-trained problem of the annotation model, which is crucial in practice but overlooked in interactive annotation.
  • Introducing Araida that involves a KNN module and an error-aware integration strategy to alleviate the under-trained problem by facilitating coordination between the two model annotators (i.e., the vanilla annotation model and KNN).
  • Demonstrating the efficacy of Araida in enhancing suggestion accuracy, reducing human corrections, and showcasing its flexibility to combine with various annotation models through extensive experiments.
2 Related Work
Our research is tied to interactive data annotation, human analogical reasoning (KNN), and retrieval-based language models. We provide a literature review and highlight our differences.
Interactive Data Annotation.
Interactive data annotation aims to reduce human annotation effort by incorporating an annotation model that suggests labels to human annotators during an interactive process (Klie et al., 2018, 2020; Le et al., 2021). The annotation model must be sample-efficient because, when we start a new annotation task, there are few labeled examples to learn from. Current studies focus on employing active learning (Klie et al., 2018; Laws et al., 2011; Casanova et al., 2020; Li et al., 2021; Huang et al., 2023) to prioritize annotating examples more likely to improve model accuracy. While active learning can reduce the required training data to some extent, it may not be effective in limited data annotation scenarios or when complex hypotheses or semantics are to be learned Dasgupta (2005); Rietz and Maedche (2021). Another approach is to employ LLMs for automatic data annotation, which have demonstrated strong performance under zero-shot and few-shot settings He et al. (2023); Gilardi et al. (2023). However, such performance might not be consistent for difficult tasks, as they may even perform worse than fine-tuned small language models Xiao et al. (2023). Regardless of whether we use active learning and whether we use a classic or LLM-based annotation model, our empirical evidence demonstrates that Araida can effectively decrease the amount of human corrections required.
KNN and Analogical Reasoning.
While KNN has been extensively utilized in NLP community (Wang et al., 2019; Liu et al., 2023; Wang et al., 2022b), its underlying mechanism is often overlooked. To shed light on this, cognitive studies (Lake et al., 2017, 2015; Mitchell, 2021) revealed that the KNN inference process aligns with human analogical reasoning, enabling efficient learning (Lake et al., 2017, 2015). In particular, analogical reasoning establishes connections between relevant aspects of the current task and past experiences, forming abstractions that enhance human reasoning capabilities (Mitchell, 2021). In this context, KNN facilitates sample-efficient learning by leveraging similarities between the example to be labeled and previously annotated examples, resulting in exemplary solutions (Bautista et al., 2016), which reduce the training data requirement.
Retrieval-Based Language Models.
There is growing interest in enhancing the output of language models by incorporating a retrieval module (usually KNN or alike) that interpolates with a datastore built from the training data Khandelwal et al. (2019); Kassner and Schütze (2020). Compared to vanilla language models, retrieval-based models ground the predictions in labeled training examples, potentially yielding better explainability and sample efficiency (Asai et al., 2023). This approach has shown promising results in tasks such as machine translation Khandelwal et al. (2021); Liu et al. (2023), named entity recognition Wang et al. (2022b), and question answering Kassner and Schütze (2020). While some studies have explored the use of dynamically adjusted combination weights between the language model and the retrieval module (Wang et al., 2021; Zheng et al., 2021; Jiang et al., 2021), our method differs significantly for two main reasons: 1) Different tasks. We are the pioneers in introducing KNN to the interactive data annotation task, whereas these methods are primarily designed for machine translation. 2) Different techniques. We adjust the weight by estimating the error of model predictions for each data point (e.g., sentence), whereas these methods learn the weight for each token without error estimation.
3 Araida: The Proposed Method
We present Araida, an analogical reasoning-based method for interactive data annotation that provides an annotation reference to the annotation model by retrieving previously human-labeled examples in the proximity of the example in consideration. We detail the KNN inference module in Section 3.1 and the error-aware integration strategy in Section 3.2. Finally, the optimization details are provided in Section 3.3.
Task Formalization and Overview.
Let $X$ denote the dataset that needs to be annotated, with $C$ being the number of classes. Given a data batch $x_{t}$ at time $t$, the annotation model $f_{t}$ predicts label vectors $f_{t}(x_{t})\in R^{|x_{t}|\times C}$, and the KNN module $g_{t}$ infers label vectors $g_{t}(x_{t})\in R^{|x_{t}|\times C}$ using previously annotated data stored in a datastore $A_{t}$. Then, we estimate the probability $\lambda_{t}\in R^{|x_{t}|\times 1}$ of the annotation model’s predictions $f_{t}(x_{t})$ being reliable, i.e., $argmax(f_{t}(x_{t}))=y_{t}$, where $argmax(\cdot)$ returns indices of the classes with the highest predicted probability and $y_{t}\in R^{|x_{t}|\times 1}$ are the ground truth labels. Finally, we use $\lambda_{t}$ to weigh the two predictions $f_{t}(x_{t})$ and $g_{t}(x_{t})$ through a linear weighted combination:
$F_{t}(x_{t})=\lambda_{t}\cdot f_{t}(x_{t})+(1-\lambda_{t})\cdot g_{t}(x_{t}).$ (1)
Notably, closed-source language models such as ChatGPT produce discrete labels rather than predicted distributions. Therefore, we cannot combine its predictions with KNN’s using linear combination. In such case, we use binary values (00 or $1$) for $\lambda_{t}$, which acts as a function allocation to determine whether $f_{t}$ or $g_{t}$ should apply to each example. Once the human approves or corrects the final suggestions, the datastore $A_{t}$ is updated with the newly arrived data batch and its corresponding labels. In addition, the annotation model $f_{t}$ (if applicable), KNN module $g_{t}$, and weighting strategy $\lambda_{t}$ are updated via back-propagation.
3.1 KNN Inference
We utilize a weighted KNN to perform inference, defined as $g_{t}(x_{t}^{i})=\frac{\sum_{a\in\rho_{i}}w_{a}y_{a}}{\sum_{a\in\rho_{i}}w_{a}}$, where $\rho_{i}\in A_{t}$ is the $k$ nearest neighbors of each example $x_{t}^{i}$, $y_{a}$ corresponds to the human annotation of each neighbor $a\in\rho_{i}$. The similarity between $x_{t}^{i}$ and $a$ is measured by $w_{a}=\frac{1}{d(x_{t}^{i},a)}$, where $d(x_{t}^{i},a)=\|w_{knn}(x_{t}^{i}-a)\|_{2}$ is a distance metric parameterized by $w_{knn}$. We use the similarity measure to retrieve $\rho_{i}$. To avoid overconfidence in KNN inference, we apply label smoothing to the labels of the retrieved neighbors. Specifically, we set $y_{a}=y_{a}(1-\alpha)+\alpha/C$, where $\alpha=1-\frac{1}{C}$.
Datastore Maintenance Strategy.
The datastore $A_{t}$ consisting of historically annotated data grows in size as the interactive annotation continues, causing the KNN’s retrieval to be less time-efficient. To address this issue, we impose a constraint on the maximum datastore size using a pre-defined hyperparameter. We propose a class-aware maintenance strategy. Precisely, if $A_{t}$ exceeds its budget, data that is from the majority class  [^2] The majority class refers to the class with the highest frequency in $A_{t}$. and most similar to its class prototype [^3] The class prototypes are the average of the feature vectors in $A_{t}$ that belong to each class. is discarded first. This strategy ensures that the datastore contains as many labeled data from different classes as possible while minimizing the impact on the class prototype. Appendix A.1 and A.3 present experiments using different datastore sizes and maintenance strategies.
3.2 Error-aware Integration Strategy
Motivation.
A popular method to combine annotations from two models (i.e., annotation model and KNN) is to use a weighted linear combination with a constant weight  [^4] Equivalent to when $\lambda_{t}(x_{t})$ in Eq.1 is a constant. (Liu et al., 2023; Wang et al., 2022b). However, assuming that one model consistently outperforms the other on all unlabeled data is unrealistic. Furthermore, both models are updated with each new batch of data in interactive data annotation, and their relative performance will alter, making it infeasible to find the optimal weight through a one-off hyperparameter tuning. To address this issue, we propose an error-aware integration strategy that automatically assigns weights to different models, relying more on KNN inference when the annotation model’s prediction is estimated to be inaccurate.
Error Estimation of Model Annotation.
We base on the intuition that if the model $f_{t}$ consistently makes mistakes on previous examples similar to the current data point $x_{t}^{i}$, then its prediction $f_{t}(x_{t}^{i})$ will likely be incorrect. To achieve this, we parameterize the integration strategy $\lambda_{t}$ as a neural network to learn from the customized input.
  • Customized Inputs. For each data point $x_{t}^{i}$, we derive the input $\mathbb{x}_{t}^{i}$ to the integration strategy $\lambda_{t}$, which considers the local error estimation $E_{t}^{i}$ and local density $D_{t}^{i}$. Specifically, $E_{t}^{i}$ is a vector with elements $e_{t,j}^{i}=\mathbb{1}[argmax(f_{t}(a_{j}))=y_{a_{j}}]$ indicates if the annotation model $f_{t}$ predicted correctly on each annotated example $a_{j}\in\rho_{i}$ in the $k$ nearest neighbors of $x_{t}^{i}$. The local density $D_{t}^{i}$ is a distance vector, with each element being $d(x_{t}^{i},a_{j})$. These two vectors are combined using the element-wise multiplication operator $\odot$ to create the input: $\mathbb{x}_{t}^{i}=D_{t}^{i}\odot E_{t}^{i}-D_{t}^{i}\odot(1-E_{t}^{i})$. Notably, $\mathbb{x}_{t}^{i}$ measures the error regularity of $x_{t}^{i}$ among its neighborhood, as the more positive values in the vector $\mathbb{x}_{t}^{i}$, the less likely $f_{t}$ would make an error on $x_{t}^{i}$.
[…]
3.3 Optimization of Araida
To simplify the optimization process, we independently optimize the annotation model $f_{t}$, KNN model $g_{t}$, and error-aware integration strategy $\lambda_{t}$. We treat human feedback $y_{t}$ as the ground truth following previous studies on interactive annotation Klie et al. (2018, 2020); Le et al. (2021). It is worth noting that Araida supports any task-specific annotation model and uses its corresponding loss function $\ell_{f}$ to update the parameters. Combining the $\ell_{d}$ loss and the negative log-likelihood loss $\ell_{g}$ to optimize KNN, we formulate the final loss function as follows:
$\begin{split}\mathcal{L}(f,g,\lambda)=\sum_{i=1}^{B_{t}}&\ell_{f}(y_{i},f_{t}(x_{i}))+\ell_{g}(y_{i},g_{t}(x_{i}))\\ &+\ell_{d}(y_{i},f_{t}(x_{i}),\lambda_{t}),\end{split}$ (2)
where $B_{t}$ represents the total data accumulated until round $t$. There are two challenges to optimizing this objective function. Firstly, the operator used in KNN to retrieve the $k$ nearest neighbors is not differentiable. To address this problem, we utilize the Gumbel-softmax-based re-parameterization trick (Jang et al., 2016) to facilitate the optimization process. Secondly, the loss function $\mathcal{L}$ presents a bi-level optimization problem, where the optimization of $\lambda_{t}$ is nested within the optimization problems of $f_{t}$ and $g_{t}$. As a result, we update $f_{t}$, $g_{t}$, and $\lambda_{t}$ iteratively using a coordinate-descent approach. Formally, at each optimization iteration $k$, we have network parameters $\theta_{f}^{k}$, $\theta_{g}^{k}$, and $\theta_{\lambda}^{k}$ corresponding to $f^{k}$, $g^{k}$, and $\lambda^{k}$. The update procedures are as follows:
$\begin{split}\theta_{f}^{k+1}&=\theta_{f}^{k}-\bigtriangledown_{f}\mathcal{L}(f,g^{k},\lambda^{k}),\\ \theta_{g}^{k+1}&=\theta_{g}^{k}-\bigtriangledown_{g}\mathcal{L}(f^{k},g,\lambda^{k}),\\ \theta_{\lambda}^{k+1}&=\theta_{\lambda}^{k}-\bigtriangledown_{\lambda}\mathcal{L}(f^{k+1},g^{k+1},\lambda).\end{split}$ (3)
4 Experiments
We conduct extensive experiments to assess Araida’s effectiveness in the limited data annotation scenario. Our primary focus is to assess whether Araida can decrease the human effort required for corrections by providing more precise annotations at various stages of the annotation process (see Section 4.2). Furthermore, we perform a comprehensive examination to investigate the behavior and impact of KNN and the error-aware integration strategy (see Section 4.3). Additional analysis of our error-aware integration strategy is presented in Section 4.4. Lastly, we analyze the sensitivity of the parameters in Appendices A.
Dataset # Val. Classes
WN18RR 3,034 Hypernym; Derivation; Member; Component; Synset; Synonym; Verb group; Instance of hypernym;
FreeBase 5,116 Contains; Country; Track_role; Profession; Group_role; Adjoins; Film_release; Nutrient
IMDB 5,000 Positive; Negative
SST-5 1,101 Strong positive; Positive; Neutral; Negative; Strong negative
Table 1: Statistics of datasets. For each dataset, we randomly sample 5K examples from the original training dataset to form the unlabeled data, and the validation dataset is taken from the original dataset. Table 4 presents the mapping from the original categories to the categories we use.
Annotation Model Without Active Learning (AL) With Active Learning (AL)
Word-level Annotation Sentence-level Annotation Word-level Annotation Sentence-level Annotation
WN18RR FreeBase IMDB SST-5 WN18RR FreeBase IMDB SST-5
Dist./FT 50.44±1.02 32.47±12.37 70.18±8.43 36.02±3.14 47.77±0.91 26.92±10.85 66.98±6.31 35.20±3.76
Dist./FT + Araida 52.16±1.37 43.02±6.43 79.33±2.81 37.21±3.03 49.54±0.84 39.06±4.57 75.84±1.14 37.02±2.50
LLaMa2 31.35±1.89 24.41±1.77 80.21±2.64 37.83±1.64 31.35±1.89 24.41±1.77 80.21±2.64 37.83±1.64
LLaMa2 + Araida 45.15±1.96 38.20±1.91 88.47±2.03 42.03±1.88 46.33±2.01 38.79±1.96 89.68±2.25 42.61±1.91
LLaMa2sft 58.24±2.79 53.11±1.38 94.06±9.02 46.84±6.30 59.28±2.57 55.39±2.01 95.13±10.15 47.45±6.17
LLaMa2sft + Araida 60.74±2.33 55.23±1.46 95.15±9.32 49.62±5.98 61.02±2.18 56.71±2.09 95.88±12.27 49.51±6.03
Table 2: Machine cumulative accuracy (MCA) scores using various methods. We report the averaged results across varying amounts of data. LLaMa2 with AL has identical performance as LLaMa2 because the vanilla LLaMa2 model is not updated during the interactive annotation process. Thus, the data order does not impact the performance. When combined with Araida, the performances w/ and w/o AL are different because the KNN and integration strategy models are updated.
4.1 Experimental Setup
Tasks & Datasets.
We experiment with word- and sentence-level annotation tasks. These tasks have been highlighted as crucial in various web applications  Yao et al. (2021); Marcos-Pablos and García-Peñalvo (2020); Lee et al. (2022). To simulate the scenario of limited data annotation, we follow Dou et al. (2019) by imposing dataset size restrictions, ranging from $1K$ to $5K$. Table 1 overviews the dataset statistics.
Evaluation Metric.
We aim to minimize the total human corrections (i.e., the total model suggestion errors) annotating a given amount of data using the interactive annotation process. Therefore, we report the Machine Cumulative Accuracy (MCA), defined as the total correct suggestions divided by the total suggestions for different dataset sizes. To assess the performance of each method in the limited data annotation scenario, we present the mean and the corresponding standard deviation by varying the dataset size ({1K, 2K, 3K, 4K, 5K}).
Annotation Models.
To verify the generalizability of Araida, we apply it in conjunction with different annotation models:
  • Classic annotation models. We utilize lightweight annotation models following previous works Desmond et al. (2021); Chen et al. (2020); Hedderich et al. (2021). Specifically, for word-level tasks, we use a distributional model Roller et al. (2014); Kober et al. (2021) with pretrained GloVe word embeddings embedding Pennington et al. (2014) [^9] We did not use contextualized embeddings because the word-level task in our experiment has no context.. For sentence-level tasks, we use FastText Joulin et al. (2017) to derive the sentence embeddings. We denote this baseline as Dist./FT.
  • LLM-based annotators. We also use large language models (LLMs) as annotation models, whose few-shot and in-context learning capabilities might help with the limited data annotation process. We experiment with LLaMa2-7B Touvron et al. (2023) and ChatGPT Ouyang et al. (2022)  [^10] The gpt-3.5-turbo checkpoint in OpenAI’s API (https://platform.openai.com/docs/models/gpt-3-5).. We use zero-shot and few-shot prompts for ChatGPT (denoted as ChatGPTzero and ChatGPTfew). Detailed prompts can be found in Table 7. For LLaMa2, we consider both the vanilla LLaMa2 that uses the same zero-shot prompts as ChatGPTzero and LLaMa2sft, which is fine-tuned using an open-source toolkit [^11] https://github.com/Alpha-VLLM/LLaMA2-Accessory during the interactive annotation process. Fine-tuning data examples can be found in Table 8.
Impact of Active Learning.
Regardless of the annotation model, the sequence of the data to annotate also affects the amount of human corrections. Intuitively, if we show unambiguous examples first, few corrections are needed. However, the annotation model and KNN may not learn to handle more challenging examples and may subsequently make more mistakes. Previous studies focused on applying active learning in interactive annotation to enhance the annotation models’ sample efficiency Laws et al. (2011); Klie et al. (2018); Li et al. (2021). To study the impact of active learning in the limited data annotation scenario, we compare an uncertainty-based active learning method with random data ordering for different annotation models with and without Araida. Note that we omit the ChatGPT with AL results because we are unable to estimate its prediction uncertainty accurately.
Implementation Details.
All experiments are carried out on a machine with Intel(R) Xeon(R) Gold 5317 CPU @ 3.00GHz and a GeForce RTX 3090 GPU. For simplicity, we implement our integration strategy using a three-layer, fully-connected network with ReLu activation and dropout. For KNN, we set $k=20$ for $\rho_{t}$. KNN runs in the embedding space of text-embedding-ada-002 Neelakantan et al. (2022) when combining with ChatGPT. For Dist./FT and LLaMa2 models, KNN runs in the corresponding model’s embedding space. Moreover, we leave the details of LLM prompts and mode fine-tuning examples in Appendix D.
Figure 3: MCA scores using ChatGPT-based methods. We omit the ChatGPT with AL results because we are unable to estimate its prediction uncertainty. Araida can further improve ChatGPT’s performance
4.2 Main Result
We evaluate the effectiveness of Araida in enhancing the model annotation quality, hence reducing human correction effort. Table 2 and Figure 3 show the machine cumulative accuracy scores averaged across varying amounts of data ($\{1K,2K,3K,4K,5K\}$) for different experimental settings. We make the following observations:
Araida reduces human corrections consistently. As illustrated in Table 2 and Figure 3, Araida consistently improves the model suggestion accuracy across different annotation models and annotation tasks. Specifically, it achieves a 13.06% performance gain in MCA on Dist./FT, 3.85% on LLaMa2sft, 16.80% on Dist./FT with AL, 2.61% on LLaMa2sft with AL, 30.48% on LLaMa2, 8.81% on ChatGPTzero, and 1.55% on ChatGPTfew. On average, it achieves an 11.02% performance gain in model annotation accuracy, translating into a reduction in human correction effort in practice.
Active learning does not always help. When comparing the annotation models with and without active learning, we observe that active learning does not always improve model performance and can even harm the performance in some instances (e.g., Dist./FT). We hypothesize that the more challenging cases selected by active learning might require more training data for the models to correct their predictions Dasgupta (2005); Rietz and Maedche (2021), which are unavailable in limited data annotation scenarios. Instead of relying on an annotation model alone, Araida acts as a posthoc “plug-in” that fixes the annotation model’s mistakes using retrieved annotation references and yields a robust improvement under various settings with and without active learning.
LLMs are strong annotation models. Araida can improve them further. LLMs perform better than classic distributional models, especially for sentence-level tasks. Furthermore, LLaMa2 with fine-tuning consistently outperforms the vanilla LLaMa2 model, and few-shot ChatGPT consistently beats its zero-shot counterpart. Interestingly, comparing these two pairs of annotation models, we observe that Araida brings a more substantial improvement to weaker models. When the annotation models are already strong (in the case of LLaMa2sft and ChatGPTfew), Araida is more conservative in making corrections, yielding a smaller but consistent improvement. It demonstrates Araida’s robustness to combine with annotations models with different performances.
Annotation Model Word-level Annotation Sentence-level Annotation
WN18RR FreeBase IMDB SST-5
Dist./FT
Araida 52.16$\pm$1.37 43.02$\pm$6.43 79.33$\pm$2.81 37.21$\pm$3.03
- w/o KNN 50.44$\pm$1.02 32.47$\pm$12.37 70.18$\pm$8.43 36.02$\pm$3.14
- w/o $f(\cdot)$ 50.28$\pm$3.01 41.52$\pm$5.87 78.48$\pm$0.68 35.02$\pm$1.37
- w/ const. 51.85$\pm$4.77 41.78$\pm$6.78 78.58$\pm$3.94 37.07$\pm$2.68
LLaMa2sft
Araida 60.74$\pm$2.33 55.23$\pm$1.46 95.15$\pm$9.32 49.62$\pm$5.98
- w/o KNN 58.24$\pm$2.79 53.11$\pm$1.38 94.06$\pm$9.02 46.84$\pm$6.30
- w/o $f(\cdot)$ 45.23$\pm$2.58 39.82$\pm$3.41 89.13$\pm$0.55 42.93$\pm$1.71
- w/ const. 58.24$\pm$2.79 53.11$\pm$1.38 94.06$\pm$9.02 46.84$\pm$6.30
ChatGPTzero
Araida 50.42$\pm$1.37 43.11$\pm$1.83 93.52$\pm$1.24 48.41$\pm$1.06
- w/o KNN 47.62$\pm$2.89 38.85$\pm$2.23 92.44$\pm$0.95 41.30$\pm$1.72
- w/o $f(\cdot)$ 48.73$\pm$2.64 41.30$\pm$5.04 90.61$\pm$0.26 45.84$\pm$1.12
- w/ const. 48.73$\pm$2.64 41.30$\pm$5.04 92.44$\pm$0.95 45.84$\pm$1.12
ChatGPTfew
Araida 52.53$\pm$1.67 52.26$\pm$3.44 94.92$\pm$1.02 55.12$\pm$1.36
- w/o KNN 51.30$\pm$1.72 51.60$\pm$3.15 94.78$\pm$0.99 53.85$\pm$2.03
- w/o $f(\cdot)$ 48.73$\pm$2.64 41.30$\pm$5.04 90.61$\pm$0.26 45.84$\pm$1.12
- w/ const. 51.30$\pm$1.72 51.60$\pm$3.15 94.78$\pm$0.99 53.85$\pm$2.03
Table 3: Ablation study on the KNN and the error-aware integration strategy modules. We report the MCA scores using various methods, averaging results with different amounts of data. Error-aware integration strategy effectively coordinates the two annotators.
Figure 4: Analyzing our integration strategy with the Dist./FT model. The solid lines show the MAC scores of the annotation model $f(\cdot)$, separated by examples with $\lambda>0.5$ (higher weights assigned to the annotation model $f(\cdot)$) and $\lambda\leq 0.5$ (higher weights assigned to KNN). The dotted line shows KNN’s performance on the latter set.
4.3 Ablation Study
This section aims to perform a comprehensive examination to investigate the behavior and impact of the KNN and out integration strategy. We conduct an ablation study to analyze the effectiveness of each component. In particular, we consider the following baselines:
  • Araida w/o KNN: Using the annotation model alone to suggest labels. Equivalent to $\lambda_{t}(\cdot)=1$.
  • Araida w/o f($\cdot$): Using KNN alone to suggest labels. Equivalent to $\lambda_{t}(\cdot)=0$.
  • Araida w/ const.: Using a constant $\lambda_{t}^{*}$ value for all examples, in contrast to Araida, which varies $\lambda_{t}$ for different examples. In our experiments, we report the result with the best $\lambda_{t}^{*}$ tuned on the validation set.
The ablation test results are presented in Table 3. Due to limited space, we omit the result for vanilla LLaMa2, which is much weaker than other LLM-based baselines. The detailed observations are provided below.
KNN is a strong stand-alone annotator. Table 3 reveals that although KNN (Araida w/o f($\cdot$)) does not match the performance of Araida, it obtains comparable accuracy as using the annotation model alone (Araida w/o KNN). This result is surprising, given KNN’s simplicity compared to the annotation models (including LLMs). Such the results also highlight that significance and effectiveness of analogical reasoning, allowing humans to reason more effectively Mitchell (2021).
Error-aware integration strategy effectively coordinates the two annotators. Table 3 shows that Araida’s error-aware integration strategy achieves consistent performance gain compared to using a constant weight $\lambda_{t}^{*}$. This in turn also confirms the original intention behind our construction of the error-aware integration strategy: it is infeasible to find the optimal weight through a one-off hyperparameter tuning. Here, it is worth noting that ChatGPT outputs discrete labels rather than probabilistic vectors. In this case, the constant strategy reduces into an indicator function: if $\lambda_{t}^{*}>0.5$, $F_{t}(x)=f_{t}(x),\forall x_{t}$; otherwise, $F_{t}(x)=g_{t}(x),\forall x_{t}$. In this case, Araida w/ const. prefers ChatGPT or KNN based on their performance. However, it cannot improve the annotation quality beyond the individual components (i.e., the KNN and the annotation model) due to the discrete output of ChatGPT.
4.4 Qualitative Analysis
To shed light on how the error-aware integration strategy works, we measure the cumulative accuracy of the annotator model and KNN throughout the annotation process for each dataset and present the result with the Dist./FT annotation model in Figure 4. We also separate the cases where $\lambda>0.5$ (higher weights assigned to the annotation model $f(\cdot)$) and $\lambda\leq 0.5$ (higher weights assigned to KNN). Our observations are as follows.
Firstly, we observe that there is a substantial gap between the two solid lines ($MCA_{f}$ on $\lambda>0.5$ and $MCA_{f}$ on $\lambda\leq 0.5$), showing that our error estimation model effectively identifies cases where $f(\cdot)$ is likely to error and assigns it a lower weight. Secondly, KNN reaches a reasonable accuracy much faster than the annotation model at the initial stage of the interactive annotation process. Even as the number of annotated examples increases, its accuracy is still higher than the annotation model $f(\cdot)$ when $\lambda\leq 0.5$. This result reveals that KNN compensates for the performance deficiencies of $f(\cdot)$ on data where it is more likely to make a mistake. Therefore, combining KNN using the error-aware integration strategy in Araida leads to an overall improvement in annotation quality, hence reducing human correction effort.
5 Conclusion
In interactive data annotation, an annotation model suggests labels to human annotators to verify. However, the annotation model is prone to errors when trained on limited labeled data. To tackle this challenge, we proposed Araida, an approach inspired by analogical reasoning, to compensate for the performance deficiencies of the annotation model and correct its mistakes using an error-aware integration strategy. Extensive experiments demonstrated that Araida is flexible to combine with different annotation models across various tasks and yields consistent improvement in label suggestion accuracy, which leads to a reduction of human correction effort. In this study, our method explores a new solution to bring more flexibility by allowing the human to design any preferred annotation model according to different annotation tasks. We are devoted to optimizing human-machine utilities by emphasizing the learning of task-specified concepts efficiently from a few human demonstrations. In future work, we plan to extend Araida to other annotation tasks and develop it as a general toolkit that can benefit the NLP community.
6 Acknowledgements
This work was supported in part by the National Natural Science Foundation of China (No. 62272330); in part by the Fundamental Research Funds for the Central Universities (No. YJ202219).
7 Limitations
Human Studies.
This work aims to reduce human correction effort in interactive data annotation. We follow previous work (Hwa, 2000; Kristjansson et al., 2004) to use the number of model suggestion errors to approximate the human correction effort needed. However, the actual effort needed depends on the particular example and the type of errors (e.g., whether it is obvious). Ideally, we would involve human annotators and measure the saving of annotation time. However, due to the large number of experimental settings, conducting human studies with each annotation model and ablation baseline was infeasible.
Error-Prone Human Annotation.
This paper treats human annotations as ground truth following previous studies in interactive data annotation (Klie et al., 2018; Le et al., 2021). However, uncertainty and inconsistency of human annotations do occur. We refer readers to the literature on handling error-prone human annotation, such as crowd-sourced data annotation (Larson et al., 2020).
Although human annotation errors are not the focus of this work, we explore Araida’s performance under synthesized label noise conditions. We consider the crowd-sourced data annotation scenario and assume that each human annotator $h_{i}$ makes mistakes with the latent probability $p_{e}^{i}\sim(0,0.3)$. We set the total number of annotators $O=10$ and sample their corresponding error probabilities $P_{e}=\{p_{e}^{1},p_{e}^{2},...,p_{e}^{O}\}$. Then, we sample $u_{i}$ from a uniform distribution $U(0,1)$ for each annotation. If $u_{i}\leq p_{e}^{i}$, $h_{i}$ assigns a randomly sampled incorrect label; otherwise, it assigns the correct one. We use majority voting of the 10 annotators to obtain the final annotations following Shirani et al. (2019).
We slightly modified Araida’s datastore maintenance strategy. When the datastore exceeds its budge, we discard the data from the majority class and most dis-similar to its class prototype instead of removing the most similar one (as in the original Araida strategy). This strategy may help remove incorrectly labeled data. The experimental result in Figure 5 shows that Araida still outperforms the baseline. The modified datastore maintenance strategy (Araida-dis) further improves the performance by a slight margin. Further research and more rigorous experiments are required to address the human annotation noise problem in interactive annotation.
Figure 5: MAC scores of various methods with synthesized label noise on the SST-5 dataset. Dist./FT is used as the annotation model. Araida-dis refers to Araida with a modified datastore maintenance strategy.
Latency.
The KNN component requires retrieving similar examples as the input data, which may limit our method’s time efficiency when the datastore size is large. To address this problem, besides the proposed datastore management strategy, we can also employ an efficient similarity search library such as FAISS  [^12] https://github.com/facebookresearch/faiss to speed up the retrieval.
Appendix A Hyperparameter Analysis
In this section, we provide a detailed hyperparameter analysis of Araida, including the datastore size $A_{t}$, the number of neighbors $k$ for $\rho_{t}$, and the datastore maintenance strategy. We show results only for the Dist./FT annotation model for brevity.
A.1 Datastore Size
We tune the size of the datastore $A_{t}$ in $\{100,500,1000,2000\}$ and evaluate the machine cumulative accuracy of Araida with or without active learning. We illustrate the results in Fig.6. A larger datastore size generally brings higher annotation performance since it allows us to maintain more data from past human-machine interactions. However, it also requires larger memory usage and causes longer latency. We found that a datastore size of 1000 is a reasonable tradeoff, which we utilize in our main experiments.
A.2 Top $k$ for $\rho_{t}$
We tune the number of neighbors $A_{t}$ in $\{5,10,20,50\}$ and evaluate the machine cumulative accuracy of Araida with or without active learning. As shown in Fig.7, $k=20$ seems to perform well for all tasks.
Refer to caption Refer to caption Refer to caption Refer to caption
(a) WN18RR (b) WN18RR + AL (c) FreeBase (d) FreeBase + AL
Refer to caption Refer to caption Refer to caption Refer to caption
(e) IMDB (f) IMDB + AL (g) SST-5 (h) SST-5 + AL
Figure 7: Machine Cumulative Accuracy of Araida with different $k$ for $\rho_{t}$.
Table 5: Results on comparing to fully fine-tuned models
Model WN18RR FreeBase IMDB SST-5
BERT(fully finetuned) 75.13$\pm$1.12 60.47$\pm$1.24 90.04$\pm$0.42 46.59$\pm$1.20
BERT+Araida 54.41$\pm$1.51 50.30$\pm$2.25 88.79$\pm$3.54 43.81$\pm$3.12
Dist./FT(fully finetuned) 72.64$\pm$1.05 58.50$\pm$1.36 84.87$\pm$0.81 42.12$\pm$1.44
Dist./FT+Araida 52.16$\pm$1.37 43.02$\pm$6.43 79.33$\pm$2.81 37.21$\pm$3.03
A.3 Datastore Maintenance Strategy
Refer to caption Refer to caption Refer to caption Refer to caption
(a) WN18RR (b) WN18RR + AL (c) FreeBase (d) FreeBase + AL
Refer to caption Refer to caption Refer to caption Refer to caption
(e) IMDB (f) IMDB + AL (g) SST-5 (h) SST-5 + AL
Figure 8: Machine Cumulative Accuracy of Araida with different datastore maintenance strategies.
We propose a class-aware datastore maintenance strategy for Araida, which removes labeled examples from the majority class most similar to their class prototype. Compared to the conventional First-In-First-Out (FIFO) strategy (Dekel et al., 2005), our method ensures that 1) the datastore contains as much data from different classes as possible, and 2) the class prototype is least affected after the removal. We compare with two variants of Araida, including Araida w/ FIFO and Araida w/ class FIFO. The former discards the oldest example regardless of the class; the latter discards the oldest example belonging to the majority class in the datastore.
As shown in Fig.8, Araida w/ FIFO can suffer from a sudden decrease in performance, as it may remove important examples arbitrarily. After integrating the class information, Araida w/ class FIFO removes the oldest analogy from the majority class, achieving a comparative performance to Araida. However, Araida still performs better when the number of classes increases (e.g., FreeBase).
Appendix B Results on Comparing to Fully Fine-tuned Model
We utilized up to 5K of data to simulate the limited data annotation task, and we updated the parameters of the annotation model as the interactive annotation process progressed. To validate the effectiveness of the fully fine-tuned model, we use the remaining data from the original dataset, excluding the 5K data and the validation data, as the training data to fully fine-tune the model. Next, we employ this model for interactive annotation. Since it is already fully fine-tuned, we do not update the model parameters during the annotation process. Due to time constraints, we are currently only considering scenarios where active learning has not been adopted. Additionally, we only fine-tune small annotation models (i.e., BERT and Dist./FT).
Based on the results in the Table 5, it is evident that the annotation performance of the fully fine-tuned model surpasses that of our method. This suggests that while our method currently offers a lightweight approach to aid data annotation, which can to some extent enhance the sample efficiency of the annotation model, there is still potential for improvement in future research.Nevertheless, it’s important to emphasize that in real-world data annotation scenarios, we typically don’t have access to fully fine-tuned models initially, as we lack labeled datasets. While one might turn to LLMs like ChatGPT, our findings in Figure 3 indicate that our method could potentially improve upon ChatGPT even further.
Appendix C Results on Smaller-sized Transformer-based Model
With a wide array of language models to choose from, we faced the challenge of not being able to test all available models. To address this, we selected three prominent models (GloVe, LLaMa2, and ChatGPT) based on our hardware resource capabilities in our main experiments. While open to conducting further experiments, our focus was limited by time constraints, leading us to concentrate solely on fine-tuning BERT (w/o AL), with the outcomes detailed in Table 6. After implementing Araida, we observed a significant enhancement in the quality of annotations by using our Araida.
Table 6: Results on smaller-sized Transformer-based model
Model WN18RR FreeBase IMDB SST-5
BERT 53.32$\pm$1.02 46.12$\pm$3.23 85.38$\pm$7.16 41.08$\pm$3.41
BERT+Araida 54.41$\pm$1.51 50.30$\pm$2.25 88.79$\pm$3.54 43.81$\pm$3.12
BERT+Araida w/o f 51.03$\pm$2.11 44.63$\pm$4.04 80.29$\pm$1.09 38.51$\pm$1.45
Appendix D Prompts and Fine-Tuning Examples for LLM-Based Annotation Models
Table 7 presents prompts used for zero-shot and few-shot annotation. Table 8 shows fine-tuning examples for LLaMa2sft.
Dataset Prompts
IMDB
input: If you like original gut wrenching laughter…it. Great Camp!!!.
output: positive
input: I saw this movie when I was about 12 when it … There are no rules.
output: negative
You need to identify the sentiments of the following sentences, output positive or negative.
{INPUTS}
SST-5
input: The gorgeously elaborate continuation …Tolkien ’s Middle-earth.
output: Strong Positive
input: Singer/composer Bryan Adams contributes a slew of…, spirit of the piece.
output: Positive
input: You’d think by now …with hearts of gold.
output: Neutral
input: This isn’t a new idea .
output: Negative
input: A sour little movie at … What was it all for ?
output: Strong Negative
Identify the sentiment of each paragraph,
you have five options: ’Strong Positive’, ’Positive’, ’Neutral’, ’Negative’ or ’Strong Negative’
{INPUTS}
WN18RR
input: ability unfitness
output: antonym
input: dissent debating
output: entailment
input: abandonment apostasy
output: hypernym
input: abandonment discard
output: hyponym
input: Afghanistan Afghan
output: member
input: abandonment abandonment
output: synonym
Identify the semantic relation of the each word pair,
you have eight options: ’component’, ’synset’,…, ’hypernym’, ’derivation’, ’member’, ’synonym’
You MUST only output the semantic relation word for each input!
{INPUTS}
FreeBase
input: Libya Egypt
output: adjoins
input: Honolulu Punahou
output: contains
input: Bobsleigh Netherlands
output: country
input: Blackbriar Lithuania
output: film_release
input: Autoharp Guitar
output: group_role
input: IceCream Water
output: nutrient
input: Shriya Actor
output: profession
input: Cello Pennywhistle
output: track_role
Identify the semantic relation of the each word pair,
you have eight options: ’contains’, ’country’, ’track_role’, ’profession’, ’group_role’, ’adjoins’, ’film_release’, ’nutrient’.
only output the semantic relation.
{INPUTS}
Table 7: Prompts for different datasets to obtain the annotation. We remove the few-shot demonstrations in the prompts in the zero-shot scenarios.
Dataset Finetuning data example for LLaMa2sft
IMDB
{"instruction": "Identify the sentiment of the following paragraph, output ’positive’ or ’negative’.",
"input": "The cast played Shakespeare.Shakespeare lost.I appreciate that this is trying to bring Shakespeare to the masses,
but why ruin something so good. Is it because ’The Scottish Play’ is my favorite Shakespeare? I do not know.
What I do know is that a certain Rev Bowdler (hence bowdlerization) tried to do something similar in the
Victorian era.In other words, you cannot improve perfection.I have no more to write but as I have to
write at least ten lines of text (and English composition was never my forte I will just have to keep
going and say that this movie, as the saying goes, just does not cut it.",
"output": "negative"},
SST-5
{"instruction": "Identify the sentiment of the following paragraph,
you have five options: ’Strong Positive’, ’Positive’, ’Neutral’, ’Negative’ or ’Strong Negative’.",
"input": "The gorgeously elaborate continuation of The Lord of the Rings trilogy is so huge that a
column of words can not adequately describe co-writer/director Peter Jackson ’s expanded vision of J.R.R.
Tolkien ’s Middle-earth .",
"output": "Strong Positive"},
WN18RR
{"instruction": "Identify the semantic relation of the following word pair,
you have eight options: ’antonym’,…, ’synonym’.",
"input": "a.m. A.M.",
"output": "synonym"},
FreeBase
{"instruction": "Identify the semantic relation of the following word pair,
you have eight options: ’contains’, …, ’nutrient’.",
"input": "Autoharp Guitar",
"output": "group_role"},
Table 8: Finetuning data examples for LLaMa2sft